# Mind the gaps: The fraught road to quantum advantage

### v1

@@ Abstract
Quantum computing is advancing rapidly, yet substantial gaps separate today’s noisy intermediate-scale quantum (NISQ) devices from tomorrow’s fault-tolerant application-scale (FASQ) machines. We identify four related hurdles along the road ahead: (i) from error mitigation to active error detection and correction, (ii) from rudimentary error correction to scalable fault tolerance, (iii) from early heuristics to mature, verifiable algorithms, and (iv) from exploratory simulators to credible advantage in quantum simulation. Targeting these transitions will accelerate progress toward broadly useful quantum computing.
@@ I Introduction
The promise of quantum computing has attracted enthusiastic attention from companies, investors, and governments, as well as scientists, fueling the progress of NISQ technology. But NISQ computers are not error corrected, and the relatively high gate error rates of today’s NISQ devices severely limit their computational power; therefore, few practical uses of this technology have emerged to date. It has long been known how to overcome this limitation [7, 8]. NISQ technology will be superseded by fault-tolerant application scale (FASQ) computers that exploit quantum error-correcting codes to run a wide variety of useful applications. Quantum hardware has reached a sufficiently advanced stage that the voyage from NISQ to FASQ is gathering momentum in companies and research laboratories. We can envision fault-tolerant quantum computers with logical gate error rates that are many orders of magnitude better than the underlying physical gate error rates of hardware platforms. But this transition, still nascent, will unfold gradually on a timescale that remains highly uncertain.
@@
Several insightful recent articles [13, 14, 15, 16, 17, 18] provide valuable perspective on the quest for useful applications of quantum technology, and Ref. [19] provides a useful overview of currently known quantum algorithms. Our article covers some of the same ground, with a particular emphasis on the four gaps along the road from from NISQ that are highlighted above. To guide the eye, we assign to each challenge a pictorial (and clearly highly subjective) “gap score,” where a larger gap indicates a more demanding task.
@@ III From protected quantum memory to scalable fault-tolerant quantum computation
Though the onset of the QEC era may be a noteworthy technological milestone, it is still in its very early stages and the pace of progress is hard to predict. In particular, while the asymptotic overhead cost of QEC scales reasonably with the size of the computation we wish to perform, in practice this cost is quite daunting from the perspective of current technology. To illustrate the predicament, consider the surface code, a particularly promising approach to early quantum fault tolerance because it can tolerate relatively strong physical noise and syndrome extraction requires only geometrically local processing in a two-dimensional layout [80, 81]. Based on numerical simulations [82], the probability of error $P_{\text{logical}}$ for each protected logical operations is related to the probability of error $p_{\text{phys}}$ for each physical two-qubit gate by
@@
Continuing progress in quantum error correction and fault tolerance will soon enable quantum technology to enter the megaquop regime in which circuits with a million or more two-qubit gates can be executed reliably [12]. A megaquop machine will be able to perform some tasks that are beyond the reach of classical NISQ, or analog quantum devices. Experience with these early fault-tolerant quantum computers will guide efforts to scale up toward more capable systems, and inform the quest for useful applications.
@@ IV From near-term quantum heuristics to mature quantum algorithms
That said, the potential value of NISQ variational algorithms remains an open issue. A positive suggestion is that the performance might be substantially improved by strategically choosing the initial parameters of the variational search (a so-called “warm start”) [126, 127, 128, 129]. For example, when seeking a low-energy state of a local Hamiltonian, one could start with an approximation provided by a classical heuristic algorithm, or when solving a combinatorial optimization problem one could start with the solution to a linear relaxation of the problem that is efficiently solvable. Conceivably, such stratagems might help to evade barren plateaus and suboptimal local minima.
Convincing arguments for quantum advantage in end-to-end variational algorithms have been elusive, but we might make progress via “proof pockets” that establish advantage for relevant subtasks. By accumulating enough such instances, a more comprehensive picture may emerge. For example, in the quantum approximate optimization algorithm (QAOA) [114], it is rigorously established that the optimal values of variational parameters concentrate [130, 131]; hence, finding near-optimal parameters for one particular problem instance can advise a warm start in solving other instances. In addition, single-round QAOA has a rigorously established advantage over classical algorithms for problems with suitable symmetries [132]. Notably, in some cases the expensive estimation of gradients can be avoided by using a dissipative procedure to optimize the cost function [133, 134]. Further incremental steps like these will continue to deepen our understanding of the power of variational quantum algorithms.
The power of end-to-end variational quantum algorithms can be clarified by “proof pockets” that characterize subtasks of the algorithms.
@@
We do not expect quantum computers to find exact solutions to worst-case instances of NP-hard problems efficiently, but they might be able to find better approximate solutions than classical computers, or to speed up the search for good approximate solutions. Variational quantum algorithms running on FASQ machines might realize that kind of advantage, but they face similar challenges to those confronting NISQ computers, such as barren plateaus and suboptimal local minima. It is known that directly simulating QAOA is classically hard [137], but also that log-depth QAOA does not have an asymptotic quantum advantage in combinatorial optimization [138] (which does not rule out a practically useful advantage at modest depth). As an existence proof, one can construct examples of asymptotic quantum advantage in finding approximate solutions to some optimization problems such as integer programming [139, 140, 141]. This is done by mapping problems solvable via Shor’s algorithm to optimization tasks, and then invoking notions of computational learning theory [139] or a variant of the theory of probabilistically checkable proofs to show that finding an approximate solution is classically hard [140]. Oracle-based subexponential quantum advantages in adiabatic optimization have also been identified [142, 143].
@@ V Toward credible quantum advantage in quantum simulation
There are efficient quantum algorithms for preparing the ground state of a local Hamiltonian on a quantum computer and then measuring observables in that state, including the total energy [176, 177]. However, these are guaranteed to work only if an input state can be prepared that has a sizeable overlap with ground state, and this is not always feasible. A general-purpose approach to preparing a good approximation to the ground state is the adiabatic method — one simulates a time-dependent Hamiltonian that interpolates between an initial non-interacting Hamiltonian whose ground state is easily prepared and a final target Hamiltonian whose ground state is desired. This method is effective if one can identify a gapped adiabatic path that connects the target ground state to the trivial ground state. However, it is likely to fail if the path crosses a first-order phase transition where the gap vanishes and the quantum state must change sharply to remain close to the instantaneous ground state along the path. Another approach is the dissipative method, in which one simulates coupling the target system to a cold thermal bath [178, 179]. This method fails if the cooled system takes a long time to thermalize, for example because its Hamiltonian has a complex energy landscape. In the cases where a reasonably good approximation to the ground state can be prepared efficiently on a quantum computer, one might question whether the ground-state problem is really classically hard [180]. For this reason, it has been challenging to argue persuasively that particular Hamiltonians of physical interest are hard to minimize classically and easy to minimize using digital quantum simulators.
@@ VI Outlook
The rapid recent advance of quantum science and technology has drawn attention not just from scientists but also from commercial ventures and government policy makers. Quantum computers with broadly useful applications are eagerly anticipated but are challenging to realize. In this article we have emphasized that to navigate the road from today’s NISQ era to tomorrow’s FASQ machines the quantum community must “mind the gaps” that block the way. We have highlighted in particular the daunting quest for scalable fault-tolerant quantum computers and the crucial hunt for practical quantum advantages over conventional information processing. We stress, though, that the portolio of useful applications is likely to expand gradually as quantum technology advances through the megaquop, gigaquop, and teraquop regimes.
@@ Acknowledgments
We are grateful to Steve Flammia for helpful and insightful comments about this article, and to many other colleagues for years of stimulating discussions about the quest for quantum advantage. The work of J. E. has been supported by the BMFTR (DAQC, MUNIQC-Atoms, QuSol, PasQuops), the Munich Quantum Valley (K-4 and K-8), the QuantERA (HQCC), the Quantum Flagship (PasQuans2, Millenion), QuantERA (HQCC), the Clusters of Excellence MATH+ and ML4Q, the DFG (CRC183), the Einstein Foundation (Einstein Research Unit on Quantum Devices), Berlin Quantum, and the ERC (DebuQC). The work of J. P. has been supported in part by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, by the U.S. Department of Energy, Office of Science, Accelerated Research in Quantum Computing, Quantum Utility through Advanced Computational Quantum Algorithms (QUACQ) and Fundamental Algorithmic Research toward Quantum Utility (FAR-Qu), and by the National Science Foundation (PHY-2317110). The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center.

### v2

@@ Abstract
Quantum computing is advancing rapidly, yet substantial gaps separate today’s noisy intermediate-scale quantum (NISQ) devices from tomorrow’s fault-tolerant application-scale quantum (FASQ) machines. We identify four related hurdles along the road ahead: (i) from error mitigation to active error detection and correction, (ii) from rudimentary error correction to scalable fault tolerance, (iii) from early heuristics to mature, verifiable algorithms, and (iv) from exploratory simulators to credible advantage in quantum simulation. Targeting these transitions will accelerate progress toward broadly useful quantum computing.
@@ I Introduction
The promise of quantum computing has attracted enthusiastic attention from companies, investors, and governments, as well as scientists, fueling the progress of NISQ technology. But NISQ computers are not error corrected, and the relatively high gate error rates of today’s NISQ devices severely limit their computational power; therefore, few practical uses of this technology have emerged to date. It has long been known how to overcome this limitation [7, 8]. NISQ technology will be superseded by fault-tolerant application-scale quantum (FASQ) computers that exploit quantum error-correcting codes to run a wide variety of useful applications. Quantum hardware has reached a sufficiently advanced stage that the voyage from NISQ to FASQ is gathering momentum in companies and research laboratories. We can envision fault-tolerant quantum computers with logical gate error rates that are many orders of magnitude better than the underlying physical gate error rates of hardware platforms. But this transition, still nascent, will unfold gradually on a timescale that remains highly uncertain.
@@
Several insightful recent articles [13, 14, 15, 16, 17, 18] provide valuable perspective on the quest for useful applications of quantum technology, and Ref. [19] provides a useful overview of currently known quantum algorithms. Our article covers some of the same ground, with a particular emphasis on the four gaps along the road from NISQ that are highlighted above. To guide the eye, we assign to each challenge a pictorial (and clearly highly subjective) “gap score,” where a larger gap indicates a more demanding task.
@@ III From protected quantum memory to scalable fault-tolerant quantum computation
Though the onset of the QEC era may be a noteworthy technological milestone, it is still in its very early stages and the pace of progress is hard to predict. In particular, while the asymptotic overhead cost of QEC scales reasonably with the size of the computation we wish to perform, in practice this cost is quite daunting from the perspective of current technology. To illustrate the predicament, consider the surface code, a particularly promising approach to early quantum fault tolerance because it can tolerate relatively strong physical noise and syndrome extraction requires only geometrically local processing in a two-dimensional layout [80, 81]. Based on numerical simulations [82, 83], the probability of error $P_{\text{logical}}$ for each protected logical operations is related to the probability of error $p_{\text{phys}}$ for each physical two-qubit gate by
@@
Continuing progress in quantum error correction and fault tolerance will soon enable quantum technology to enter the megaquop regime in which circuits with a million or more two-qubit gates can be executed reliably [12]. A megaquop machine will be able to perform some tasks that are beyond the reach of classical, NISQ, or analog quantum devices. Experience with these early fault-tolerant quantum computers will guide efforts to scale up toward more capable systems, and inform the quest for useful applications.
@@ IV From near-term quantum heuristics to mature quantum algorithms
That said, the potential value of NISQ variational algorithms remains an open issue. A positive suggestion is that the performance might be substantially improved by strategically choosing the initial parameters of the variational search (a so-called “warm start”) [127, 128, 129, 130]. For example, when seeking a low-energy state of a local Hamiltonian, one could start with an approximation provided by a classical heuristic algorithm, or when solving a combinatorial optimization problem one could start with the solution to a linear relaxation of the problem that is efficiently solvable. Classical computations might also guide the choice of the initial quantum state input to the quantum algorithm, or the manner in which the quantum algorithm seeks nearby states that achieve a better value of the cost function [130]. Conceivably, such stratagems might help to evade barren plateaus and suboptimal local minima. Conversely, the output of a quantum algorithm might inform a warm start for a classical combinatorial optimization solver [131].
Convincing arguments for quantum advantage in end-to-end variational algorithms have been elusive, but we might make progress via “proof pockets” that establish advantage for relevant subtasks or for instructive special cases. By accumulating enough such instances, a more comprehensive picture may emerge. For example, in the quantum approximate optimization algorithm (QAOA) [115], it is rigorously established that the optimal values of variational parameters concentrate [132, 133]; hence, finding near-optimal parameters for one particular problem instance can advise a warm start in solving other instances. In addition, single-round QAOA has a rigorously established advantage over classical algorithms for problems with suitable symmetries [134]. Another encouraging rigorous result identifies a family of graphs for which polynomial-time QAOA can find cuts through the graph with a cut fraction (ratio of the number of cut edges to the total number of edges) exceeding what is known to be achievable by any subexponential-time classical algorithm [135]. Notably, in some cases the expensive estimation of gradients can be avoided by using a dissipative procedure to optimize the cost function [136, 137]. Further incremental steps like these will continue to deepen our understanding of the power of variational quantum algorithms.
The power of end-to-end variational quantum algorithms can be clarified by “proof pockets” that characterize subtasks of the algorithms or establish quantum advantage for instructive special cases.
@@
We do not expect quantum computers to find exact solutions to worst-case instances of NP-hard problems efficiently, but they might be able to find better approximate solutions than classical computers, or to speed up the search for good approximate solutions. Variational quantum algorithms running on FASQ machines might realize that kind of advantage, but they face similar challenges to those confronting NISQ computers, such as barren plateaus and suboptimal local minima. It is known that directly simulating QAOA is classically hard [140], but also that log-depth QAOA does not have an asymptotic quantum advantage for a large class of sparse combinatorial optimization problems [141, 142] (which does not rule out a practically useful advantage at modest depth). As an existence proof, one can construct examples of asymptotic quantum advantage in finding approximate solutions to some optimization problems such as integer programming [143, 144, 145]. This is done by mapping problems solvable via Shor’s algorithm to optimization tasks, and then invoking notions of computational learning theory [143] or a variant of the theory of probabilistically checkable proofs to show that finding an approximate solution is classically hard [144]. Oracle-based subexponential quantum advantages in adiabatic optimization have also been identified [146, 147].
@@ V Toward credible quantum advantage in quantum simulation
There are efficient quantum algorithms for preparing the ground state of a local Hamiltonian on a quantum computer and then measuring observables in that state, including the total energy [180, 181]. However, these are guaranteed to work only if an input state can be prepared that has a sizeable overlap with the ground state, and this is not always feasible. A general-purpose approach to preparing a good approximation to the ground state is the adiabatic method — one simulates a time-dependent Hamiltonian that interpolates between an initial non-interacting Hamiltonian whose ground state is easily prepared and a final target Hamiltonian whose ground state is desired. This method is effective if one can identify a gapped adiabatic path that connects the target ground state to the trivial ground state. However, it is likely to fail if the path crosses a first-order phase transition where the gap vanishes and the quantum state must change sharply to remain close to the instantaneous ground state along the path. Another approach is the dissipative method, in which one simulates coupling the target system to a cold thermal bath [182, 183]. This method fails if the cooled system takes a long time to thermalize, for example because its Hamiltonian has a complex energy landscape. In the cases where a reasonably good approximation to the ground state can be prepared efficiently on a quantum computer, one might question whether the ground-state problem is really classically hard [184]. For this reason, it has been challenging to argue persuasively that particular Hamiltonians of physical interest are hard to minimize classically and easy to minimize using digital quantum simulators.
@@ VI Outlook
The rapid recent advance of quantum science and technology has drawn attention not just from scientists but also from commercial ventures and government policy makers. Quantum computers with broadly useful applications are eagerly anticipated but are challenging to realize. In this article we have emphasized that to navigate the road from today’s NISQ era to tomorrow’s FASQ machines the quantum community must “mind the gaps” that block the way. We have highlighted in particular the daunting quest for scalable fault-tolerant quantum computers and the crucial hunt for practical quantum advantages over conventional information processing. We stress, though, that the portfolio of useful applications is likely to expand gradually as quantum technology advances through the megaquop, gigaquop, and teraquop regimes.
@@ Acknowledgments
We are grateful to Steve T. Flammia for helpful and insightful comments about this article, and to many other colleagues for years of stimulating discussions about the quest for quantum advantage. We also thank Ieva Cepaite, Edward Farhi, Mark Goh, Sam Gutmann, Peter Love, Shaghayegh Moradirad, and Zoltan Zimboras for valuable comments. The work of J. E. has been supported by the BMFTR (DAQC, MUNIQC-Atoms, QuSol, PasQuops), the Munich Quantum Valley (K-4 and K-8), the QuantERA (HQCC), the Quantum Flagship (PasQuans2, Millenion), QuantERA (HQCC), the Clusters of Excellence MATH+ and ML4Q, the DFG (CRC183), the Einstein Foundation (Einstein Research Unit on Quantum Devices), Berlin Quantum, and the ERC (DebuQC). The work of J. P. has been supported in part by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, by the U.S. Department of Energy, Office of Science, Accelerated Research in Quantum Computing, Quantum Utility through Advanced Computational Quantum Algorithms (QUACQ) and Fundamental Algorithmic Research toward Quantum Utility (FAR-Qu), and by the National Science Foundation (PHY-2317110). The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center.